OpenStudy (anonymous):

In a certain pentagon, the interior angles are a,b,c,d,e degrees where a,b,c,d,e are integers strictly less than 180. ("Strictly less than 180" means they are "less than and not equal to" 180.) If the median of the interior angles is 61 degrees and there is only one mode, then what are the degree measures of all five angles?

OpenStudy (welshfella):

i imagine there is no one answer to this question The following list of angles would meet the requirements 59,60,61,61,149,150 ( = 540 degrees )

OpenStudy (welshfella):

ooops thats wrong - I have 6 angles there 1 only 5 in a pentagon!

OpenStudy (loser66):

and the median is not 61 as well

OpenStudy (welshfella):

61 is middle angle since its the median

OpenStudy (loser66):

i don't get what "there is only one mod" mean

OpenStudy (welshfella):

so we have -- -- 61 -- --

OpenStudy (loser66):

yes, you are right

OpenStudy (welshfella):

- i guess one angle occurs twice? or it could occur 3 times?? - thats still only one mod right?

OpenStudy (loser66):

yes, you are right

OpenStudy (welshfella):

what if we had 61 , 61,61 how many degrees left? 540 - 183 = 357 so other 2 angles could be 179 and 178

OpenStudy (welshfella):

that would fit

OpenStudy (welshfella):

61,61,61,178,179

OpenStudy (loser66):

I am pretty sure that the sentence "there is only one mod" has some meaning and it limits the answer. @satellite73

OpenStudy (loser66):

@mathmate @amistre64

OpenStudy (loser66):

@Michele_Laino

OpenStudy (welshfella):

- hasn't that list only got one mode?

OpenStudy (welshfella):

its an odd question

OpenStudy (welshfella):

- i think this might be the only solution because 3 61's is the only combination that will all ow the last 2 numbers to less than 180 degrees and different

OpenStudy (loser66):

@Kainui

OpenStudy (anonymous):

Please 61 is the Median and yes two angles can be 61 which means the other 3 have to be same to be a mode of which there can be only one. Now 61*2=122 So, 540-122= 418 418/3 = 139.33333.. Now something isn't right here or my understanding incorrect

OpenStudy (loser66):

We can turn it into one mode. :) as what welshfella said, it can be 61-61- 61 and n, n 540-3*61 = 357, that means n = 178.5 to turn it to multiple of 61, we do 61 * factor = 178.5 , hence factor = 178.5/ 61=(357/122) hence the angles are 61,61,61, (357/122)61, (357/122)61 hahaha... I am crazy.!!!

OpenStudy (anonymous):

Loser.. as the name suggest... I am serious here.. need to solve this real quick

OpenStudy (loser66):

Yes, I am a big Loser. I want to study also. That was why I tagged many others to get help for YOU. If my solution doesn't satisfy you, I am so sorry.

OpenStudy (anonymous):

u r not a loser.. I cant understand why we convert median to mode.. there is a catch here... also I tired to solve.. what is wrong in that can you suggest. Name or no name.. whats in it ? A guide can be a mis-guide.. sorry I didn't meant to be rude

OpenStudy (welshfella):

what is wrong with my solution? It fits all the requirements